# Quantum-inspired probability metrics define a complete, universal space for statistical learning

### v1

@@ A.1 Theorem 1
(Theorem 1, SI Appendix, A.1) Assume we have an affine map $T$ that is a closed embedding (a homeomorphism from the weak topology on $\mathcal{P}_{\mathrm{w}}(X)$ to its norm-closed image in $E$); its inverse $T^{-1}$ is also affine. Since $\mathcal{P}_{\mathrm{w}}(X)$ is convex, its affine image is convex. Using Lemma 1, choose a net $\mu_{\alpha}$ that is weak-* convergent to a shift-invariant generalized measure $\mu$. By Lemma 2, the image of this net $T(\mu_{\alpha})$ converges weak-* to $T(\mu)\in E^{**}$. Since $E$ is reflexive, weak-* convergence is equivalent to weak convergence, and $E^{**}\cong E$, so $T(\mu_{\alpha})$ converges weakly to $T(\mu)\in E$. Define $v_{\alpha}\equiv T(\mu_{\alpha})$ as its image in $E$, and $v\equiv T(\mu)$ as its weak limit.
@@
Recall that the Kernel Mean Embedding is the affine map $T:\mathcal{P}_{\mathrm{w}}(X)\to\mathcal{H}$ defined by $T(\delta_{x})=K_{x}$ for the kernel function $K_{x}(\,\cdot\,)\equiv K(\,\cdot\,,x)\in\mathcal{H}$ [2]. By Lemma 2, this map can be extended to all generalized measures, and by the argument in Theorem 1 (SI Appendix, A.1), given any weak-* convergent net $\mu_{\alpha}\to\mu$, the image of the limit $T(\mu)$ is contained in the norm closure of the image of $\mathcal{P}_{\mathrm{w}}(X)$.
@@ A.2 Theorem 2
Although this RKHS may look like $L^{2}(\mathbb{C}^{n})$, it only contains functions of the form $\exp(-|\mathbf{z}|^{2}/2)\overline{F(\mathbf{z})}$, with $F$ entire. This is a closed subspace of $L^{2}(\mathbb{C}^{n})$: briefly, $L^{2}$ convergence of Gaussian-weighted entire functions implies uniform convergence on compact sets, which yields a Gaussian-weighted entire limit by Morera’s theorem. The expansion of the inner product in coherent states suggests the well-known resolution of the identity operator:
@@ A.3 Theorem 3
which is now viewed as a measure on the closed set $\mathcal{C}\subset\mathrm{extr}\,\mathcal{S}$ (recall that pure states are the extreme points of the convex quantum state space $\mathcal{S}$). Measures on $\mathcal{C}$ inherit the weak topology from measures $X$, and vice versa, via the homeomorphism $\phi$. The barycenter map $T(\mu)$ is continuous from $\mathcal{P}_{\mathrm{w}}(\mathcal{C})$ to $\mathcal{S}$ with the trace norm [6], it is convex (affine), and its image is contained in the closed convex hull of $\mathcal{C}$, denoted $\overline{\mathrm{co}}(\mathcal{C})$.
@@
where $\langle m|$ represents the embedded dual vector corresponding to the measure $m$. (Embedding as dual vector $\langle m|\in\mathcal{H}^{*}$ instead of as a vector $|m\rangle$ is conventional because it allows the expected value of a function $|f\rangle$ in the RKHS to be calculated as $\langle m|f\rangle$.)
@@
fails to be injective, i.e. for distinct measures $m_{1},m_{2}$ we have $T(m_{1})=T(m_{2})$. Applying the operator $T(m_{1})-T(m_{2})$ to the witness vector $|w\rangle$ gives:
@@ A.5 Theorem 5
For the multivariate case, if a function $f(\mathbf{x})$ on $\mathbb{R}^{n}$ has a compactly-supported Fourier transform, it extends to an entire function $f(\mathbf{z})$ of exponential type on $\mathbb{C}^{n}$, i.e., there are positive constants $A,B$ with $|f(\mathbf{z})|\leq Ae^{B|\mathbf{z}|}$. Such a function is of exponential type in each complex coordinate alone, treating the other coordinates as fixed. That is, if $f(\mathbf{z})$ is of exponential type, then for $z_{k}$ alone we have:
@@
for a constant $\tau_{2}$. Iterating in this fashion proves the lemma. With some additional effort (not needed for this proof), one can show that the constants $\tau_{j}$ define the support of the distributional Fourier transform as the box $\tau_{1},\tau_{1}]\times[-\tau_{2},\tau_{2}]\times\ldots\times[-\tau_{n},\tau_{n}]$. ∎
@@
Recall from eq. 7 of Lemma 4 that the Q-representation of the operator $\hat{F}$ has the following absolutely convergent power series:
@@
$Q_{F}(\mathbf{x}+i\mathbf{y})=e^{-(|\mathbf{x}|^{2}+|\mathbf{y}|^{2})}\!\sum_{\mathbf{n},\mathbf{m}=\mathbf{0}}^{\mathbf{\infty}}\frac{\langle\mathbf{n}|\hat{F}|\mathbf{m}\rangle}{\sqrt{\mathbf{n}!\mathbf{m}!}}(\mathbf{x}-i\mathbf{y})^{\mathbf{n}}(\mathbf{x}+i\mathbf{y})^{\mathbf{m}}$
Written in this form, it is evident that the right-hand side is a real analytic function of $\mathbf{x},\mathbf{y}$ that extends to an entire function for complex $\mathbf{x},\mathbf{y}$. But the left-hand side cannot be extended to complex arguments, because the expression $\mathbf{x}+i\mathbf{y}$ for complex $\mathbf{x},\mathbf{y}$ does not capture the $4n$ real degrees of freedom contained in the two complex vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n}$.
@@
$\frac{\langle\overline{\mathbf{x}}+i\overline{\mathbf{y}}|\hat{F}|\mathbf{x}+i\mathbf{y}\rangle}{\langle\overline{\mathbf{x}}+i\overline{\mathbf{y}}|\mathbf{x}+i\mathbf{y}\rangle}$ $=e^{-(|\mathbf{x}|^{2}+|\mathbf{y}|^{2})}\!\sum_{\mathbf{n},\mathbf{m}}\!\frac{\langle\mathbf{n}|\hat{F}|\mathbf{m}\rangle}{\sqrt{\mathbf{n}!\mathbf{m}!}}(\mathbf{x}\!-\!i\mathbf{y})^{\mathbf{n}}(\mathbf{x}\!+\!i\mathbf{y})^{\mathbf{m}}$
@@
if and only if the Fock-basis matrix elements of $\hat{F}$ are equal to the power-series coefficients, $\langle n|\hat{F}|m\rangle=F_{nm}$; since $F(x,y)$ is real, these coefficients must be Hermitian, $F_{nm}=\overline{F_{mn}}$. Thus, we define the operator $\hat{F}$ to have these matrix elements. At this stage, it is not at all clear that they define a bounded operator. Nonetheless, by construction, this operator will satisfy:
@@
where the first line expresses the trace in terms of coherent states, the second line inserts the resolution of the identity, and the third line identifies the kernel $K_{f}(w,z)=\langle w|\hat{f}|z\rangle$. The operator $\hat{f}$ is Hilbert-Schmidt (and defines a function in the Gaussian RKHS) if and only if this kernel is square-integrable over all $w$ and $z$. In contrast, the Schur test requires only that the kernel be absolutely integrable over $z$ for each fixed $w$, and vice versa—a much weaker requirement. As a result, the IPM witness functions for the QPM are a much broader class of functions compared with the RKHS functions used in MMD. For instance, while all functions in the Gaussian RKHS must vanish at infinity, we have shown that all real, bounded, band-limited functions on $\mathbb{R}^{n}$ are dual functions in the QPM defined by the Fock embedding.
@@ C Replacing MMD with a QPM
and define the MMD as the Hilbert space distance between the embedded vectors (this gives the biased MMD, i.e., V-statistic):
@@
Writing the difference vector $|D\rangle=|\mathbb{P}\rangle-|\mathbb{Q}\rangle$ as:
@@
    Where MMD was calculated from the Gram matrix $\langle x_{i}|x_{j}\rangle=K(x_{i},x_{j})$, calculate a new Gram matrix using the square root of the kernel, i.e. $\langle x_{i}|x_{j}\rangle=\sqrt{K(x_{i},x_{j})}$. One must ensure here that $\sqrt{K}$ is also a positive definite kernel.
@@ D QPMs in Generative Moment Matching Networks (GMMNs)
To generate the images in Figure 1, we trained for 200 epochs on the MNIST training set (50,000 images) using batches of 1000 and the PyTorch AdamW optimizer with a learning rate of 0.005 and default parameters. (See https://github.com/logansmccarty/QPM for results and code.) A single training loop trained two models together with the only difference being the objective function (QPM versus MMD). The MMD results are qualitatively similar to those from [10], with many visible artifacts that do not improve beyond 200 epochs. In contrast, the images generated using QPM appear much more similar to the MNIST samples.

### v2

@@ Abstract
Acknowledgments
The AI model “GPT-5 Thinking” from OpenAI [13] assisted with checking proofs in the SI Appendix.
@@ A.1 Theorem 1
(Theorem 1, SI Appendix, A.1) Assume we have an affine closed embedding map $T$ (a homeomorphism from the weak topology on $\mathcal{P}_{\mathrm{w}}(X)$ to its norm-closed image in $E$); its inverse $T^{-1}$ is also affine. Since $\mathcal{P}_{\mathrm{w}}(X)$ is convex, its image is convex. Using Lemma 1, choose a net $\mu_{\alpha}\in\mathcal{P}_{\mathrm{w}}(X)$ that is weak-* convergent to a shift-invariant generalized measure $\mu$. By Lemma 2, the image of this net $T(\mu_{\alpha})$ converges weak-* to $T(\mu)\in E^{**}$. Since $E$ is reflexive, weak-* convergence is equivalent to weak convergence, and $E^{**}\cong E$, so $T(\mu_{\alpha})$ converges weakly to $T(\mu)\in E$. Define $v_{\alpha}\equiv T(\mu_{\alpha})$ as its image in $E$, and $v\equiv T(\mu)$ as its weak limit.
@@
Recall that the Kernel Mean Embedding is the affine map $T:\mathcal{P}_{\mathrm{w}}(X)\to\mathcal{H}$ defined by $T(\delta_{x})=K_{x}$ for the kernel function $K_{x}(\,\cdot\,)\equiv K(\,\cdot\,,x)\in\mathcal{H}$ [2]. By Lemma 2, this map can be extended to all generalized measures, and by the argument in Theorem 1 (SI Appendix, A.1), given any weak-* convergent net of probability measures $\mu_{\alpha}\to\mu$, the image of the limit $T(\mu)$ is contained in the norm closure of the image of $\mathcal{P}_{\mathrm{w}}(X)$.
@@ A.2 Theorem 2
Although this RKHS may look like $L^{2}(\mathbb{C}^{n})$, it only contains functions of the form $\exp(-|\mathbf{z}|^{2}/2)\overline{F(\mathbf{z})}$, with $F$ entire. This is a closed subspace of $L^{2}(\mathbb{C}^{n})$ [4]: briefly, $L^{2}$ convergence of Gaussian-weighted entire functions implies uniform convergence on compact sets, which yields a Gaussian-weighted entire limit by Morera’s theorem. The expansion of the inner product in coherent states suggests the well-known resolution of the identity operator:
@@ A.3 Theorem 3
which is now viewed as a measure on the closed set $\mathcal{C}\subset\mathrm{extr}\,\mathcal{S}$ (recall that pure states are the extreme points of the convex quantum state space $\mathcal{S}$). Measures on $\mathcal{C}$ inherit the weak topology from measures on $X$, and vice versa, via the homeomorphism $\phi$. The barycenter map $T(\mu)$ is continuous from $\mathcal{P}_{\mathrm{w}}(\mathcal{C})$ to $\mathcal{S}$ with the trace norm [6], it is convex (affine), and its image is contained in the closed convex hull of $\mathcal{C}$, denoted $\overline{\mathrm{co}}(\mathcal{C})$.
@@
where $\langle m|$ represents the embedded dual vector corresponding to the measure $m$. Embedding as dual vector $\langle m|\in\mathcal{H}^{*}$ instead of as a vector $|m\rangle$ is conventional because it allows the expected value of a function $|f\rangle$ in the RKHS to be calculated as $\langle m|f\rangle$.
@@
fails to be injective, i.e. for some distinct measures $m_{1},m_{2}$ we have $T(m_{1})=T(m_{2})$. Following the argument above and applying the operator $T(m_{1})-T(m_{2})$ to the witness vector $|w\rangle$ gives:
@@ A.5 Theorem 5
For the multivariate case, if a function $f(\mathbf{x})$ on $\mathbb{R}^{n}$ has a compactly-supported distributional Fourier transform, it extends to an entire function $f(\mathbf{z})$ of exponential type on $\mathbb{C}^{n}$, i.e., there are positive constants $A,B$ with $|f(\mathbf{z})|\leq Ae^{B|\mathbf{z}|}$. Such a function is of exponential type in each complex coordinate alone, treating the other coordinates as fixed. That is, if $f(\mathbf{z})$ is of exponential type, then for $z_{k}$ alone we have:
@@
for some constant $\tau_{2}$. Iterating in this fashion proves the lemma. With additional effort (not needed for this proof), one can show that the constants $\tau_{j}$ define the support of the distributional Fourier transform as the box $\tau_{1},\tau_{1}]\times[-\tau_{2},\tau_{2}]\times\ldots\times[-\tau_{n},\tau_{n}]$. ∎
@@
Recall from eq. 7 of Lemma 4 that the Q-representation of the operator $\hat{F}$ has the absolutely convergent power series:
@@
$Q_{F}(\mathbf{x}+i\mathbf{y})=e^{-(\mathbf{x}^{2}+\mathbf{y}^{2})}\!\sum_{\mathbf{n},\mathbf{m}=\mathbf{0}}^{\mathbf{\infty}}\frac{\langle\mathbf{n}|\hat{F}|\mathbf{m}\rangle}{\sqrt{\mathbf{n}!\mathbf{m}!}}(\mathbf{x}-i\mathbf{y})^{\mathbf{n}}(\mathbf{x}+i\mathbf{y})^{\mathbf{m}}$
(we notate $\mathbf{x}^{2}=\mathbf{x}\cdot\mathbf{x}=x_{1}^{2}+\ldots+x_{n}^{2}$, and same for $\mathbf{y}^{2}$; these are holomorphic squares not norm-squares.) Written in this form, it is evident that the right-hand side is a real analytic function of $\mathbf{x},\mathbf{y}$ that extends to an entire function for complex $\mathbf{x},\mathbf{y}$. But the left-hand side cannot be extended to complex arguments, because the expression $\mathbf{x}+i\mathbf{y}$ for complex $\mathbf{x},\mathbf{y}$ does not capture the $4n$ real degrees of freedom contained in the two complex vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n}$.
@@
$\frac{\langle\overline{\mathbf{x}}+i\overline{\mathbf{y}}|\hat{F}|\mathbf{x}+i\mathbf{y}\rangle}{\langle\overline{\mathbf{x}}+i\overline{\mathbf{y}}|\mathbf{x}+i\mathbf{y}\rangle}$ $=e^{-(\mathbf{x}^{2}+\mathbf{y}^{2})}\sum_{\mathbf{n},\mathbf{m}}\!\frac{\langle\mathbf{n}|\hat{F}|\mathbf{m}\rangle}{\sqrt{\mathbf{n}!\mathbf{m}!}}(\mathbf{x}\!-\!i\mathbf{y})^{\mathbf{n}}(\mathbf{x}\!+\!i\mathbf{y})^{\mathbf{m}}$
@@
if and only if the Fock-basis matrix elements of $\hat{F}$ are equal to the power-series coefficients, $\langle n|\hat{F}|m\rangle=F_{nm}$; since $F(x,y)$ is real, these coefficients must be Hermitian, $F_{nm}=\overline{F_{mn}}$. Thus, we define the operator $\hat{F}$ to have these matrix elements. At this stage, it is not at all clear that they define a bounded operator. Nonetheless, $\hat{F}$ with these specified Fock-basis matrix elements will satisfy:
@@
where the first line expresses the trace in terms of coherent states, the second inserts the resolution of the identity, and the third identifies the kernel $K_{f}(w,z)=\langle w|\hat{f}|z\rangle$. The operator $\hat{f}$ is Hilbert-Schmidt (and defines a function in the Gaussian RKHS) if and only if this kernel is square-integrable over all $w$ and $z$. In contrast, the Schur test requires only that the kernel be absolutely integrable over $z$ for each fixed $w$, and vice versa—a much weaker requirement. As a result, even on a compact space, the IPM witness functions for the QPM are a much broader class of functions compared with the RKHS functions used in MMD. For instance, while all functions in the Gaussian RKHS must vanish at infinity, we have shown that all real, bounded, band-limited functions on $\mathbb{R}^{n}$ are dual functions in the QPM defined by the Fock embedding.
@@ C Replacing MMD with a QPM
so MMD is the Hilbert space distance between these vectors:
@@
(If $(\mathbb{P},\mathbb{Q})$ are sampled i.i.d. from $(\mu,\nu)$, this is a biased estimate of $\text{MMD}(\mu,\nu)$, i.e. V-statistic.) Writing the difference vector $|D\rangle=|\mathbb{P}\rangle-|\mathbb{Q}\rangle$ as:
@@
    Where MMD was calculated from the Gram matrix $\langle x_{i}|x_{j}\rangle=K(x_{i},x_{j})$, calculate a new Gram matrix using the square root of the kernel, i.e. $\langle x_{i}|x_{j}\rangle=\sqrt{K(x_{i},x_{j})}$. One must ensure here that $\sqrt{K}$ is also a valid (positive definite) kernel.
@@ D QPMs in Generative Moment Matching Networks (GMMNs)
To generate the images in Figure 1, we trained for 200 epochs on the MNIST training set (50,000 images) using batches of 1000 and the PyTorch AdamW optimizer with a learning rate of 0.005 and default parameters. (See https://doi.org/10.5281/zenodo.17024967 for results and code.) A single training loop trained two models together with the only difference being the objective function (QPM versus MMD). The MMD results are qualitatively similar to those from [10], with many visible artifacts that do not improve beyond 200 epochs. In contrast, the images generated using QPM appear much more similar to the MNIST samples.